Truth Hartman is married, claims four allowances and earns $1,416.00 semimonthly. Compute Hartman's tax using the percentage method. a.$76.00 b.$22.40 c.$94.72 d.$34.68
Answer #1

Answer is option C

C. $22.40

Taxable earnings for federal income tax withheld = 1416 – (175*4) = $716

Income tax withheld= (716-492)*10% = $22.40